OpenStudy (anonymous):

i swear im right ! Which of the follwoing is equivalent to 5 singma 3n +2 ? 3 i did 3(3)+2 +3(4)+2 +3(5)+2 and got 42. my teacher says this is incorrect !

ganeshie8 (ganeshie8):

pls post the full q... n varies from wat to wat ?

OpenStudy (anonymous):

n=3 sorry

OpenStudy (unklerhaukus):

\[\sum\limits_{n=3}^53n+2\]\[=(3(3)+2) +(3(4)+2) +(3(5)+2)\]\[=(9+2)+(12+2)+(15+2)\]\[=9+12+15+3\times2\]\[=21+15+6\]\[=21+21\]\[=42\]

OpenStudy (anonymous):

\[\sum_3^5(3n+2)\]?

OpenStudy (unklerhaukus):

is the two is definitely in the sum ?

OpenStudy (anonymous):

unkle thats what i did !

OpenStudy (unklerhaukus):

maybe there is another method that the teacher is looking for,

OpenStudy (anonymous):

i dont know any other methods we didnt learn any other methods. ...

OpenStudy (anonymous):

don't give it another thought, you have three numbers to add and you did it correctly either the teacher is wrong, (it happens) or there is a typo somewhere

OpenStudy (unklerhaukus):

\[\sum\limits_{n=3}^5(3n+2)\]\[=3\sum\limits_{n=3}^5n+\sum\limits_{n=3}^52\]\[=3(3+4+5)+(2+2+2)\]\[=3(12)+6\]\[=36+6\]\[=42\]

OpenStudy (mathmate):

...unless the "+2" at the end does not belong to the summation if what you posted is part of a longer expression.